the length of a rectangle is greater than its width by 4 cm. if the perimeter of the rectangle is 80 ' FIND THE SIDES OF THE RECTANGLE.

Found 2 solutions by JulietG, stanbon:
Answer by JulietG(1812)

The perimeter is a fence around. It requires 2 lengths of fence + 2 widths of fence. (P = 2L + 2W)
You're given the following information:
P = 80
L = W + 4
Substitute the value for L in the original equation:
P (80) = 2(W+4) + 2W
Distribute
80 = 2W + 8 + 2W
Add the Ws
80 = 4W + 8
Subtract 8 from each side
72 = 4W
Divide each side by 4
18 = W
.
Since you know the width is 18, you can find the length by adding 4. 18+4=22
Let's make certain by putting in the values:
2(22) + 2(18) = 80
44 + 36 does indeed equal 80.

Answer by stanbon(75887)

the length of a rectangle is greater than its width by 4 cm. if the perimeter of the rectangle is 80 ' FIND THE SIDES OF THE RECTANGLE.
--------------------
P = 2(L + W)
80 = 2(W+4 + W)
-------------------
40 = 2W+4
2W = 36
-------
Width = 18 cm
-----------------------
Length = W+4 = 22 cm
==========================
Cheers,
Stan H.
==============